OpenStudy (anonymous):

with sides of x + 2 and an equilateral triangle of sides with length 2x have equal perimeters.

OpenStudy (anonymous):

1 2 4

OpenStudy (anonymous):

I think your question's missing something. Is it a square of sides x+2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright, so perimeter of the square = 4(x+2) and perimeter of the equilateral triangle = 2x*3 4(x+2) = 2x*3 4x + 8 = 6x 8 = 2x x = 4